OpenStudy (anonymous):

If 2n+1 is an odd number, what are the next two consecutive odd numbers?

OpenStudy (anonymous):

2n+3 and 2n+5. Try it out my good sir. if n=1, 2n+1=3, 2n+3=5, 2n+5, 7 3,5,7.

OpenStudy (anonymous):

lol thanks my good sir. I appreciate it.